3x-8y=-21 x=2-1/3y

Answer by macston(5194)

x=2-1/3y
3x-8y=-21 Substitute for x.
3(2-(1/3)y)-8y=-21
6-y-8y=-21
-9y=-27
y=3 ANSWER 1: y=3
x=2-1/3y
x=2-(1/3)3
x=2-1=1 ANSWER 2: x=1
CHECK:
3x-8y=-21
3(1)-8(3)=-21
3-24=-21
-21=-21